help me to answr this this ;
2x+5y=7
3x-2y=1

Answer by rapaljer(4671)

2x+5y=7
3x-2y=1

To make the y coefficients subtract out, multiply both sides of the first equation times 2, and the second equation times 5:

2(2x+5y=7)
5(3x-2y=1)

4x+10y=14
15x-10y=5

Add these equations together:
19x=19
x=1

Substitute x=1 into the first equation:
2x+5y=7
2*1+5y=7
2+5y=7
5y=5
y=1

Substitute both values back into the second equation to check:
3x-2y=1
3*1-2*1=1
3-2=1 It checks!!

For more detailed explanations, examples, and exercises, including solutions in LIVING COLOR on my website by clicking on my tutor name "rapaljer" anywhere in algebra.com. Look for the pages on my home page, either "Basic, Intermediate, and College Algebra" or "Math in Living Color". In either of these pages, select "Basic Algebra" and then "Chapter 4." Look for the appropriate sections, for an easier to understand explanation than traditional textbooks!!

R^2